# Bosonic and fermionic coherence of N-partite states in the background of a dilaton black hole

### v1

@@ II The quantization of field in GHS dilaton spacetime
$ds^{2}=-\bigg{(}\frac{r-2M}{r-2D}\bigg{)}dt^{2}+\bigg{(}\frac{r-2M}{r-2D}\bigg{)}^{-1}dr^{2}+r(r-2D)d\Omega^{2},$ (1)
where $D$ and $M$ are parameters related to the dilaton field and the mass of the black hole, respectively. The relationship among $D$, the charge $Q$, and $M$ is described by the equation $D=Q^{2}/2M$. In this paper, we consider $G=c=\hbar=\kappa_{B}=1$.
@@ II.1 Bosonic field
where $Y_{lm}(\theta,\varphi)$ represents a scalar spherical harmonic on the unit two-sphere and $h(r)=\sqrt{r(r-2D)}$. We can easily obtain the radial equation
@@ III.1 N-partite coherences of GHZ and W states for bosonic field
$W^{B}_{123\ldots N+n}$ $=$ $\frac{1}{\sqrt{N}}\bigg{\{}(1-e^{-8\pi\omega(M-D)})^{\frac{n}{2}}|1\rangle_{{}_{N}}|0\rangle_{{}_{N-1}}\cdots|0\rangle_{{}_{n+1}}\big{[}\bigotimes^{n}_{i=1}$ (31)
@@
$+$ $(1-e^{-8\pi\omega(M-D)})^{\frac{n}{2}}|0\rangle_{{}_{N}}|1\rangle_{{}_{N-1}}\cdots|0\rangle_{{}_{n+1}}\big{[}\bigotimes^{n}_{i=1}$
@@
$\times$ $|0\rangle_{{}_{n+1}}\big{[}\bigotimes^{n}_{i=2}\alpha_{i}|m_{i}\rangle_{out}|m_{i}\rangle_{in}\big{]}$
$\times$ $\big{[}\sum^{\infty}_{m_{1}=0}e^{-4\pi m_{1}\omega(M-D)}\sqrt{m+1_{1}}|m+1_{1}\rangle_{out}|m_{1}\rangle_{in}\big{]}\bigg{\}}.$
@@
Figure 1: Quantum coherences $C^{B}(GHZ)$ and $C^{B}(W)$ of bosonic field as a function of the dilaton $D$ for different $N$, $p$, $n$, where we have fixed $M=\omega=1$.
Fig.1 shows how the dilaton $D$ of the black hole influences the coherences of GHZ and W states for bosonic field. From Fig.1(a) and (b), we can observe the physically accessible coherence $(p=n)$ of bosonic field monotonically decreases to a fixed value with the increase of the dilaton $D$, which means that N-partite coherence can survive in the dilaton black hole, while bosonic steering experiences catastrophic behavior, such as ”sudden death” P18 . For the fixed $N$, the physically accessible coherence decreases monotonically with the increase of the number of bosons $n$ near the event horizon of the dilaton black hole. In Fig.1(c) and (d), we also observe that the inaccessible coherence $(p=18,n=20)$ initially increases to a maximum before decreasing to a fixed value as the dilaton $D$ increases. However, the inaccessible coherence $(p=8,n=20)$ monotonically increases with increasing dilaton $D$ in Fig.1(e) and (f). In other words, the physically inaccessible N-partite coherence increases either monotonically or non-monotonically with the increase of the dilaton $D$, depending on the relative numbers between the accessible and the inaccessible modes. In addition, the coherence of W state is always bigger than that of GHZ state.
Figure 2: Quantum coherences of GHZ and W states for bosonic field as functions of $p$ and $n$ for fixed $M=\omega=1$ and $D\longrightarrow 1$.
@@ III.2 N-partite coherences of GHZ and W states for fermionic field
Figure 3: Quantum coherences $C^{F}(GHZ)$ and $C^{F}(W)$ for fermionic field as a function of the dilaton $D$ for different $N$, $p$, $n$, where we have fixed $M=\omega=1$.
Fig.3 shows how the dilaton $D$ of the black hole influences the coherences of GHZ and W states for fermionic field. From Fig.3, we can see that the curve of the fermionic field is similar to the curve of the bosonic field in curved spacetime. Through the analysis, we get the conclusions of fermionic coherence: (i) when $p=n$, the N-partite physically accessible coherence first decreases and then reaches to a fixed value with the dilaton $D$; (ii) the inaccessible coherence increases either monotonically or non-monotonically with the dilaton $D$, depending on the relative numbers between the accessible and the inaccessible modes; (iii) the fermionic coherence of W state is always larger than that of GHZ state.
Figure 4: Quantum coherences of GHZ and W states for fermionic field as functions of $p$ and $n$ for fixed $M=\omega=1$ and $D\longrightarrow 1$.
@@
Therefore, the above analytical expression reflects that the fermionic coherence of GHZ state remains unchanged with the increase of $p$ in the limit of $D\rightarrow M$. In this case, the coherence of GHZ state for fermionic field can be written in a simpler form
@@ III.3 Compared coherence with entanglement of bosonic and fermionic fields
In previous papers, these results have shown that fermionic entanglement is larger than bosonic entanglement in relativistic frame, which means that fermionic entanglement is better suitable for handling relativistic quantum information tasks L30 ; L31 ; L32 ; L33 ; L34 ; L35 ; L36 ; L37 . Generally, bosonic entanglement disappears in the case of an extreme black hole, while fermionic entanglement can survive indefinitely L33 ; P31 . Naturally, a question arose: is fermionic coherence larger than bosonic coherence in dilaton spacetime? It is essential to select appropriate quantum resources and types of particles to effectively handle relativistic quantum information tasks in dilaton spacetime. In this section, we will compare coherence with entanglement of bosonic and fermionic fields in curve spacetime.
@@
Figure 5: Quantum entanglements and coherences of tripartite states for bosonic and fermionic fields as a function of the dilaton $D$ for fixed $M=\omega=1$.
Fig.5 shows how the dilaton $D$ of the black hole affects quantum entanglements and coherences of GHZ and W states for bosonic and fermionic fields. From Fig.5(a) and (b), we can see that for GHZ or W state, fermionic entanglement is always larger than bosonic entanglement. Comparing Fig.5(a) with (b), we find that quantum entanglement of GHZ state is larger than quantum entanglement of W state in dialton spacetime. Unlike quantum entanglement in dilaton spacetime, Fig.5(c) and (d) show that for GHZ or W state, bosonic coherence is larger than fermionic coherence. However, Fig.5(c) and (d) also show that the coherence of W state is larger than the coherence of GHZ state. Through the above analysis, the properties of quantum entanglements and coherences of GHZ and W states for bosonic and fermionic fields are completely opposite. Hence, some conclusions in the dilaton black hole can be obtained as follow: (i) fermionic entanglement is more suitable for processing relativistic quantum information compared to bosonic entanglement; (ii) bosonic coherence is more effective for processing relativistic quantum information than fermionic coherence; (iii) quantum entanglement of GHZ state is a superior quantum resource compared to quantum coherence of W state (iv) the coherence of W state is a better quantum resource than the coherence of GHZ state. In other words, for different types of particles, we should choose appropriate quantum resources to deal with quantum information tasks.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (Grant Nos. 12205133), LJKQZ20222315, and the Special Fund for Basic Scientific Research of Provincial Universities in Liaoning under grant NO. LS2024Q002.

### v2

@@ II The quantization of field in GHS dilaton spacetime
$ds^{2}=-\bigg{(}\frac{r-2M}{r-2\widetilde{\mathcal{D}}M}\bigg{)}dt^{2}+\bigg{(}\frac{r-2M}{r-2\widetilde{\mathcal{D}}M}\bigg{)}^{-1}dr^{2}+r(r-2\widetilde{\mathcal{D}}M)d\Omega^{2},$ (1)
where $\widetilde{\mathcal{D}}$ and $M$ are parameters related to the dilaton field and the mass of the black hole, respectively. The relationship among $\widetilde{\mathcal{D}}$, the charge $Q$, and $M$ is described by the equation $\widetilde{\mathcal{D}}=Q^{2}/2M^{2}$. In this paper, we consider $G=c=\hbar=\kappa_{B}=1$.
@@ II.1 Bosonic field
where $Y_{lm}(\theta,\varphi)$ represents a scalar spherical harmonic on the unit two-sphere and $h(r)=\sqrt{r(r-2D)}$ with $D=Q^{2}/2M$. We can easily obtain the radial equation
@@ III.1 N-partite coherences of GHZ and W states for bosonic field
$W^{B}_{123\ldots N+n}$ $=$ $\frac{1}{\sqrt{N}}\bigg{\{}(1-e^{-8\pi\omega(M-D)})^{\frac{n}{2}}|1\rangle_{{}_{N}}|0\rangle_{{}_{N-1}}\cdots|0\rangle_{{}_{n+1}}\big{[}\sum^{\infty}_{m_{i}=0}\bigotimes^{n}_{i=1}$ (31)
@@
$+$ $(1-e^{-8\pi\omega(M-D)})^{\frac{n}{2}}|0\rangle_{{}_{N}}|1\rangle_{{}_{N-1}}\cdots|0\rangle_{{}_{n+1}}\big{[}\sum^{\infty}_{m_{i}=0}\bigotimes^{n}_{i=1}$
@@
$\times$ $|0\rangle_{{}_{n+1}}\big{[}\sum^{\infty}_{m_{i}=0}\bigotimes^{n}_{i=2}\alpha_{i}|m_{i}\rangle_{out}|m_{i}\rangle_{in}\big{]}$
$\times$ $\big{[}\sum^{\infty}_{m_{1}=0}\beta_{m_{1}}|m+1_{1}\rangle_{out}|m_{1}\rangle_{in}\big{]}\bigg{\}}.$
@@
Figure 1: Quantum coherences $C^{B}(GHZ)$ and $C^{B}(W)$ of bosonic field as a function of the dilaton $\widetilde{\mathcal{D}}$ for different $N$, $p$, $n$, where we have fixed $M=\omega=1$.
Fig.1 shows how the dilaton $\widetilde{\mathcal{D}}$ of the black hole influences the coherences of GHZ and W states for bosonic field. From Fig.1(a) and (b), we can observe the physically accessible coherence $(p=n)$ of bosonic field monotonically decreases to a fixed value with the increase of the dilaton $\widetilde{\mathcal{D}}$, which means that N-partite coherence can survive in the dilaton black hole, while bosonic steering experiences catastrophic behavior, such as ”sudden death” P18 . For the fixed $N$, the physically accessible coherence decreases monotonically with the increase of the number of bosons $n$ near the event horizon of the dilaton black hole. In Fig.1(c) and (d), we also observe that the inaccessible coherence $(p=18,n=20)$ initially increases to a maximum before decreasing to a fixed value as the dilaton $\widetilde{\mathcal{D}}$ increases. However, the inaccessible coherence $(p=8,n=20)$ monotonically increases with increasing dilaton $\widetilde{\mathcal{D}}$ in Fig.1(e) and (f). In other words, the physically inaccessible N-partite coherence increases either monotonically or non-monotonically with the increase of the dilaton $\widetilde{\mathcal{D}}$, depending on the relative numbers between the accessible and the inaccessible modes. In addition, the coherence of W state is always bigger than that of GHZ state.
Figure 2: Quantum coherences of GHZ and W states for bosonic field as functions of $p$ and $n$ for fixed $M=\omega=1$ and $\widetilde{\mathcal{D}}\longrightarrow 1$.
@@ III.2 N-partite coherences of GHZ and W states for fermionic field
Figure 3: Quantum coherences $C^{F}(GHZ)$ and $C^{F}(W)$ for fermionic field as a function of the dilaton $\widetilde{\mathcal{D}}$ for different $N$, $p$, $n$, where we have fixed $M=\omega=1$.
Fig.3 shows how the dilaton $\widetilde{\mathcal{D}}$ of the black hole influences the coherences of GHZ and W states for fermionic field. From Fig.3, we can see that the curve of the fermionic field is similar to the curve of the bosonic field in curved spacetime. Through the analysis, we get the conclusions of fermionic coherence: (i) when $p=n$, the N-partite physically accessible coherence first decreases and then reaches to a fixed value with the dilaton $\widetilde{\mathcal{D}}$; (ii) the inaccessible coherence increases either monotonically or non-monotonically with the dilaton $\widetilde{\mathcal{D}}$, depending on the relative numbers between the accessible and the inaccessible modes; (iii) the fermionic coherence of W state is always larger than that of GHZ state.
Figure 4: Quantum coherences of GHZ and W states for fermionic field as functions of $p$ and $n$ for fixed $M=\omega=1$ and $\widetilde{\mathcal{D}}\longrightarrow 1$.
@@
Therefore, the above analytical expression reflects that the fermionic coherence of GHZ state remains unchanged with the increase of $p$ in the limit of $\widetilde{\mathcal{D}}\rightarrow M$. In this case, the coherence of GHZ state for fermionic field can be written in a simpler form
@@ III.3 Compared coherence with entanglement of bosonic and fermionic fields
In previous papers, these results have shown that fermionic entanglement is larger than bosonic entanglement in relativistic frame, which means that fermionic entanglement is better suitable for handling relativistic quantum information tasks L30 ; L31 ; L32 ; L33 ; L34 ; L35 ; L36 ; L37 . Generally, bosonic entanglement disappears in the case of an extreme black hole, while fermionic entanglement can survive indefinitely L33 ; P31 ; L237 ; L238 ; L239 ; gh1 . Naturally, a question arose: is fermionic coherence larger than bosonic coherence in dilaton spacetime? It is essential to select appropriate quantum resources and types of particles to effectively handle relativistic quantum information tasks in dilaton spacetime. In this section, we will compare coherence with entanglement of bosonic and fermionic fields in curve spacetime.
@@
Figure 5: Quantum entanglements and coherences of tripartite states for bosonic and fermionic fields as a function of the dilaton $\widetilde{\mathcal{D}}$ for fixed $M=\omega=1$.
Fig.5 shows how the dilaton $\widetilde{\mathcal{D}}$ of the black hole affects quantum entanglements and coherences of GHZ and W states for bosonic and fermionic fields. From Fig.5(a) and (b), we can see that for GHZ or W state, fermionic entanglement is always larger than bosonic entanglement. Comparing Fig.5(a) with (b), we find that quantum entanglement of GHZ state is larger than quantum entanglement of W state in dialton spacetime. Unlike quantum entanglement in dilaton spacetime, Fig.5(c) and (d) show that for GHZ or W state, bosonic coherence is larger than fermionic coherence. However, Fig.5(c) and (d) also show that the coherence of W state is larger than the coherence of GHZ state. Through the above analysis, the properties of quantum entanglements and coherences of GHZ and W states for bosonic and fermionic fields are completely opposite. Hence, some conclusions in the dilaton black hole can be obtained as follow: (i) fermionic entanglement is more suitable for processing relativistic quantum information compared to bosonic entanglement; (ii) bosonic coherence is more effective for processing relativistic quantum information than fermionic coherence; (iii) quantum entanglement of GHZ state is a superior quantum resource compared to quantum coherence of W state (iv) the coherence of W state is a better quantum resource than the coherence of GHZ state. In other words, for different types of particles, we should choose appropriate quantum resources to deal with quantum information tasks.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (Grant Nos. 12205133), LJKQZ20222315, and the Special Fund for Basic Scientific Research of Provincial Universities in Liaoning under grant NO. LS2024Q002. The authors would like to thank Wentao Liu for helpful discussions.